OpenStudy (anonymous):

line 1 passes through the points (1,2) and (5,-9). Line 2 passes through (1,-5) adn (14,-4) . Are the lines parrallel, perpendicular, or neither? How do you know?

OpenStudy (mertsj):

Slope of first line = -11/4. Slope of second line = 1/13. Since the slopes are not the same nor are they negative reciprocals, the answer is neither.

OpenStudy (aroub):

line 1 ; -9-2/5-1 -11/4 line 2 : -4+5/ 14-1 1/13 The slopes are different so neither :)

OpenStudy (anonymous):

thank you